# Chain-of-Influence: Tracing Interdependencies Across Time and Features in Clinical Predictive Modelings

### v2

@@ Abstract
Chronic Kidney Disease (CKD) affects approximately 8%-16% of the global population and represents one of the most significant challenges in modern healthcare, with its gradual and irreversible progression toward End-Stage Renal Disease (ESRD) imposing substantial clinical and economic burdens NKF_CKD_2024; NCHS2019Mortality. The complexity of CKD progression stems from multifaceted interactions between clinical, demographic, and socioeconomic factors that evolve over extended time periods, making accurate prediction and effective management particularly challenging lin2013progression; li2024enhancing.
Healthcare data capturing this progression is inherently temporal and multi-dimensional, with clinical outcomes resulting from complex interactions between patient characteristics, laboratory values, medications, and comorbidities. However, accurate prediction alone is insufficient for clinical deployment—clinicians require transparent and interpretable models to confidently incorporate predictive insights into patient care caruana2015intelligible; rudin2019stop; tonekaboni2019clinicians. This need for interpretability has driven significant evolution in explainable artificial intelligence (XAI), from traditional post-hoc methods like LIME ribeiro2016should and SHAP lundberg2017unified toward more sophisticated attention-based approaches that provide intrinsic interpretability choi2016retain; choi2017gram; bardhan2024icu.
Despite these advances, current approaches to clinical prediction face fundamental limitations. Traditional methods either aggregate temporal information into static features, losing crucial temporal dynamics, or model temporal patterns while treating features independently, missing critical inter-feature relationships. Even sophisticated attention-based models like RETAIN choi2016retain, while providing interpretability through temporal and feature-level attention, fail to explicitly model how features influence each other across different time points—a critical gap in understanding disease progression pathways.
@@ 2 Related Work
Attention mechanisms marked a turning point for interpretability in clinical deep learning. The foundational model in this area, RETAIN, introduced a two-level attention mechanism that operates in reverse chronological order to mimic clinical reasoning choi2016retain. By first weighting the importance of entire patient visits and then individual clinical codes within those visits, RETAIN provides clear, intuitive explanations for its predictions. Following this, models like Dipole extended the concept by using bidirectional LSTMs to capture a more holistic view of the patient’s history, exploring various attention formulations to better understand inter-visit dependencies ma2017dipole. While these pioneering models successfully identify what clinical events are important and when they occurred, they model these events largely as independent contributors to the final prediction. They do not explicitly capture the influence pathways—how a variable at one point in time may directly affect a different variable later on.
@@
A significant challenge in clinical data is its inherent irregularity and the varying importance of events over time. To address this, a second wave of research developed more sophisticated temporal attention models. For instance, ATTAIN integrated time-decay factors into its attention weights, allowing the model to dynamically prioritize events based on their temporal proximity to the prediction time zhang2019attain. Other models introduced domain-specific structural priors. StageNet, for example, models disease progression through distinct stages, using a combination of recurrent networks and attention to learn stage-specific representations and their transitions over time gao2020stagenet. Similarly, AdaCare introduces an adaptive attention mechanism that modulates feature importance based on the patient’s evolving disease severity, captured by a separate deep learning component ma2020adacare. These approaches refine the temporal and contextual aspects of attention, but the inter-feature dependencies remain encapsulated within the black-box hidden states of recurrent networks, rather than being an explicit and interpretable output of the model.
@@
The Transformer and its self-attention mechanism provide a powerful way to model global, all-to-all interactions within a sequence vaswani2017attention. In the clinical domain, BEHRT treats a patient record as a sequence of concepts and learns contextualized representations li2020behrt; PAVE builds on this with multi-layer self-attention to identify important events and aggregate them into higher-level clinical patterns steinberg2021language; and XTSFormer introduces multi-scale temporal attention to separate short-term acute signals from long-term chronic trends zhang2024xtsformer.
@@ 3.1.2 Temporal and Feature-Level Attention
Inspired by the two-level attention framework in RETAIN choi2016retain, we employ two separate bidirectional LSTM modules that operate in parallel on $\mathbf{E}_{\text{t}}$ to extract complementary temporal and feature-level importance scores. Unlike RETAIN’s use of GRU units, our bidirectional LSTMs capture long-range dependencies from both past and future contexts, which is critical for detecting subtle temporal-feature interactions in clinical sequences.
@@ 3.1.3 Cross-Feature Transformer Layers
The feature-weighted embeddings $\widetilde{\mathbf{E}}$ are passed through $L$ stacked Transformer layers to model temporal dependencies and inter-feature interactions. Each layer $l$ applies multi-head self-attention ($N_{h}$ heads) and a feed-forward network, with residual connections and Dynamic Tanh (DyT) normalization zhu2025transformers after each sub-layer.
@@ 4.1.2 MIMIC-IV Dataset
To assess generalizability in fast-changing conditions, we use MIMIC-IV v3.1 johnson2023mimic and construct a cohort from first ICU stays (65,366 patients out of 94,458 stays). Raw ICU measurements are recorded at irregular times; for modeling, we aggregate events into 1-hour windows over the first 48 hours and retain irregularity signals via masks and time-gap features (time since last observation) across 15 variables (vitals, labs, demographics). The outcome is in-hospital mortality (10.8% prevalence).
@@ 4.3 Evaluation Metrics
We evaluate model performance using standard classification metrics. The primary metric is the Area Under the Receiver Operating Characteristic curve (AUROC) hanley1982meaning; fawcett2006introduction, which measures the model’s discriminative ability across all classification thresholds. We also report the F1-Score van1979information, the harmonic mean of precision and recall that provides balanced performance assessment for imbalanced datasets. Additionally, we assess overall prediction accuracy sokolova2009systematic, which captures the proportion of correctly classified instances across both classes. Finally, we examine precision van1979information, which measures positive predictive value and is particularly important in healthcare applications to minimize false positive predictions powers2011evaluation.
@@ 4.4 Training and Implementation Details
Before training, we impute missing values using a MICE-based technique raghunathan2001multivariate; van2007multiple applied separately on each dataset. Given the substantial class imbalance in both cohorts, we evaluate several mitigation strategies—including SMOTE, ADASYN, class-weighted loss, and Temporal SMOTE (TSMOTE)—and find that TSMOTE offers the best F1–AUROC trade-off (Detailed result is provided in Appendix C.2), so we adopt it for all main experiments. Hyperparameters are selected via grid search over model-specific search spaces (Appendix B). All models are optimized with Adam and a numerically stable BCEWithLogitsLoss, with early stopping based on validation AUROC. We report test performance as the mean and 95% confidence interval over five random seeds. Additional training details are provided in Appendix B.6.
@@ 5.2.1 Attention Pattern Comparison
At the feature level (Figure 2(b)2(c)), both models consistently identify key predictors, with diabetes emerging as the dominant risk factor, in line with clinical evidence burckhardt2017multi; li2025towards. Laboratory markers (S4, S5), healthcare expenditures (net_exp_P/O), and hemoglobin levels also rank among the top features. However, these static rankings obscure the temporal dynamics that CoI’s influence chains reveal. For example, CoI ranks eGFR as 6th versus 14th in RETAIN, but neither ranking captures how eGFR’s influence propagates through time to drive utilization patterns, or how diabetes chains through cardiovascular complications to accelerate kidney decline over specific horizons. The Chain-of-Influence paradigm addresses this gap by mapping when features become influential and how their effects cascade through clinical pathways, providing actionable insights beyond static scores.

### v3

@@ Abstract
Chronic Kidney Disease (CKD) affects approximately 8%-16% of the global population and represents one of the most significant challenges in modern healthcare, with its gradual and irreversible progression toward End-Stage Renal Disease (ESRD) imposing substantial clinical and economic burdens National Kidney Foundation (2024); Kochanek et al. (2020). The complexity of CKD progression stems from multifaceted interactions between clinical, demographic, and socioeconomic factors that evolve over extended time periods, making accurate prediction and effective management particularly challenging Lin et al. (2013); Li and Padman (2026).
Healthcare data capturing this progression is inherently temporal and multi-dimensional, with clinical outcomes resulting from complex interactions between patient characteristics, laboratory values, medications, and comorbidities. However, accurate prediction alone is insufficient for clinical deployment—clinicians require transparent and interpretable models to confidently incorporate predictive insights into patient care Caruana et al. (2015); Rudin (2019); Tonekaboni et al. (2019). This need for interpretability has driven significant evolution in explainable artificial intelligence (XAI), from traditional post-hoc methods like LIME Ribeiro et al. (2016) and SHAP Lundberg and Lee (2017) toward more sophisticated attention-based approaches that provide intrinsic interpretability Choi et al. (2016; 2017); Ma et al. (2017).
Despite these advances, current approaches to clinical prediction face fundamental limitations. Traditional methods either aggregate temporal information into static features, losing crucial temporal dynamics, or model temporal patterns while treating features independently, missing critical inter-feature relationships. Even sophisticated attention-based models like RETAIN Choi et al. (2016), while providing interpretability through temporal and feature-level attention, fail to explicitly model how features influence each other across different time points—a critical gap in understanding disease progression pathways.
@@ 2 Related Work
Attention mechanisms marked a turning point for interpretability in clinical deep learning. The foundational model in this area, RETAIN, introduced a two-level attention mechanism that operates in reverse chronological order to mimic clinical reasoning Choi et al. (2016). By first weighting the importance of entire patient visits and then individual clinical codes within those visits, RETAIN provides clear, intuitive explanations for its predictions. Following this, models like Dipole extended the concept by using bidirectional LSTMs to capture a more holistic view of the patient’s history, exploring various attention formulations to better understand inter-visit dependencies Ma et al. (2017). While these pioneering models successfully identify what clinical events are important and when they occurred, they model these events largely as independent contributors to the final prediction. They do not explicitly capture the influence pathways—how a variable at one point in time may directly affect a different variable later on.
@@
A significant challenge in clinical data is its inherent irregularity and the varying importance of events over time. To address this, a second wave of research developed more sophisticated temporal attention models. For instance, ATTAIN integrated time-decay factors into its attention weights, allowing the model to dynamically prioritize events based on their temporal proximity to the prediction time Zhang (2019). Other models introduced domain-specific structural priors. StageNet, for example, models disease progression through distinct stages, using a combination of recurrent networks and attention to learn stage-specific representations and their transitions over time Gao et al. (2020). Similarly, AdaCare introduces an adaptive attention mechanism that modulates feature importance based on the patient’s evolving disease severity, captured by a separate deep learning component Ma et al. (2020). These approaches refine the temporal and contextual aspects of attention, but the inter-feature dependencies remain encapsulated within the black-box hidden states of recurrent networks, rather than being an explicit and interpretable output of the model.
@@
The Transformer and its self-attention mechanism provide a powerful way to model global, all-to-all interactions within a sequence Vaswani et al. (2017). In the clinical domain, BEHRT treats a patient record as a sequence of concepts and learns contextualized representations Li et al. (2020); PAVE builds on this with multi-layer self-attention to identify important events and aggregate them into higher-level clinical patterns Steinberg et al. (2021); and XTSFormer introduces multi-scale temporal attention to separate short-term acute signals from long-term chronic trends Xiao et al. (2025).
@@ 3.1.2 Temporal and Feature-Level Attention
Inspired by the two-level attention framework in RETAIN Choi et al. (2016), we employ two separate bidirectional LSTM modules that operate in parallel on $\mathbf{E}_{\text{t}}$ to extract complementary temporal and feature-level importance scores. Unlike RETAIN’s use of GRU units, our bidirectional LSTMs capture long-range dependencies from both past and future contexts, which is critical for detecting subtle temporal-feature interactions in clinical sequences.
@@ 3.1.3 Cross-Feature Transformer Layers
The feature-weighted embeddings $\widetilde{\mathbf{E}}$ are passed through $L$ stacked Transformer layers to model temporal dependencies and inter-feature interactions. Each layer $l$ applies multi-head self-attention ($N_{h}$ heads) and a feed-forward network, with residual connections and Dynamic Tanh (DyT) normalization Zhu et al. (2025) after each sub-layer.
@@ 4.1.2 MIMIC-IV Dataset
To assess generalizability in fast-changing conditions, we use MIMIC-IV v3.1 Johnson et al. (2023) and construct a cohort from first ICU stays (65,366 patients out of 94,458 stays). Raw ICU measurements are recorded at irregular times; for modeling, we aggregate events into 1-hour windows over the first 48 hours and retain irregularity signals via masks and time-gap features (time since last observation) across 15 variables (vitals, labs, demographics). The outcome is in-hospital mortality (10.8% prevalence).
@@ 4.3 Evaluation Metrics
We evaluate model performance using standard classification metrics. The primary metric is the Area Under the Receiver Operating Characteristic curve (AUROC) Hanley and McNeil (1982); Fawcett (2006), which measures the model’s discriminative ability across all classification thresholds. We also report the F1-Score Powers (2020), the harmonic mean of precision and recall that provides balanced performance assessment for imbalanced datasets. Additionally, we assess overall prediction accuracy Sokolova and Lapalme (2009), which captures the proportion of correctly classified instances across both classes. Finally, we examine precision Powers (2020), which measures positive predictive value and is particularly important in healthcare applications to minimize false positive predictions Powers (2020).
@@ 4.4 Training and Implementation Details
Before training, we impute missing values using a MICE-based technique Raghunathan et al. (2001); Van Buuren (2007) applied separately on each dataset. Given the substantial class imbalance in both cohorts, we evaluate several mitigation strategies—including SMOTE, ADASYN, class-weighted loss, and Temporal SMOTE (TSMOTE)—and find that TSMOTE offers the best F1–AUROC trade-off (Detailed result is provided in Appendix C.2), so we adopt it for all main experiments. Hyperparameters are selected via grid search over model-specific search spaces (Appendix B). All models are optimized with Adam and a numerically stable BCEWithLogitsLoss, with early stopping based on validation AUROC. We report test performance as the mean and 95% confidence interval over five random seeds. Additional training details are provided in Appendix B.6.
@@ 5.2.1 Attention Pattern Comparison
At the feature level (Figure 2(b)2(c)), both models consistently identify key predictors, with diabetes emerging as the dominant risk factor, in line with clinical evidence Burckhardt et al. (2017); Li et al. (2025). Laboratory markers (S4, S5), healthcare expenditures (net_exp_P/O), and hemoglobin levels also rank among the top features. However, these static rankings obscure the temporal dynamics that CoI’s influence chains reveal. For example, CoI ranks eGFR as 6th versus 14th in RETAIN, but neither ranking captures how eGFR’s influence propagates through time to drive utilization patterns, or how diabetes chains through cardiovascular complications to accelerate kidney decline over specific horizons. The Chain-of-Influence paradigm addresses this gap by mapping when features become influential and how their effects cascade through clinical pathways, providing actionable insights beyond static scores.
@@ Appendix E Chain-of-Influence Visualizations
Figure 6: Top-55 patient-specific Chain-of-Influence graph for a single patient, showing the most influential feature–time pairs and their directed paths into the prediction node. Time Interval: $t_{0}\rightarrow prediction$.
Figure 7: Path-centric view anchored at CKD Stage 4 at $t_{3}$, highlighting the high-influence chain through subsequent visits that drives the final ESRD risk.